How can I change the subject of the formula in 1.1 y=mx+q



1.2.1 m


1.2.2 q


1.2.3 x

1
Expert's answer
2022-03-07T17:33:04-0500

1.2.1 "y=mx+q"

Subtracting q from both sides

"mx=y-q"

Dividing both sides by x

"m=\dfrac{y-q}{x}"


1.2.2 "y=mx+q"

Subtracting mx from both sides

"q=y-mx"


1.2.3 "y=mx+q"

Subtracting q from both sides

"mx=y\u2212q"

Dividing both sides by m

"x=\dfrac{y-q}{m}"